### v1

# Asymptotically good CSS-T codes exist

Abstract
We give a new construction of binary quantum codes that enables the generation of a CSS-T code from any given CSS code. Using this construction, we prove the existence of asymptotically good binary CSS-T codes, resolving a previously open problem in the literature. Furthermore, we demonstrate that the same result holds for binary quantum low-density parity check CSS-T codes, and establish the existence of asymptotically good CSS codes that support any given $Z$ rotation transversally. Finally, we analyze the structure of the logical operators corresponding to certain non-Clifford gates supported by the quantum codes obtained from our construction.
Keywords— quantum code, CSS code, CSS-T code, asymptotically good codes, LDPC code, transversal gate
Introduction
Quantum error correction is a well-known and widely used technique for protecting quantum information from corruption by noise and, consequently, enabling the possibility of constructing fault-tolerant quantum computers. An $[\![n,k,d]\!]$ binary quantum error-correcting code encodes $k$ logical qubits of information into $n$ physical qubits and can correct up to $\lfloor\frac{d-1}{2}\rfloor$ errors. By itself, a quantum error correction code acts as a memory. However, operations over the logical level are required in order to obtain fault-tolerant quantum computers capable of realizing tasks that may exhibit quantum advantage. In this sense, logical operations are realized by manipulating the physical qubits encoding the information so that the desired logical operator is applied to the encoded logical qubits. Importantly, the operations applied to the physical qubits should preserve the code space, otherwise, the way in which the logical qubit is protected would essentially be lost. Furthermore, those logical operators must be induced in a fault-tolerant manner, implying that the overall quantum computation on the logical level is reliable enough.
The most natural approach to fault tolerance is the design of transversal gates, which are formed by the tensor product of single-qubit gates. A key feature of transversal operators is that single errors do not propagate within the code blocks during their operation, making errors trackable.
In general, the Eastin–Knill theorem states that no universal gate set can be implemented on the logical level in a transversal manner for binary quantum error-correcting codes with $d>1$ [15]. Recently, a way to circumvent such a no-go theorem has been proposed under certain restrictions on the error model, but its application seems to be restricted to $d$-dimensional quantum systems [20]. There exist several universal quantum gate sets, with the Clifford+$T$ gate set being the most studied one [22]. Quantum circuits composed purely of Clifford gates are efficiently classically simulable as a result of the Gottesman–Knill theorem [18]. Implementing only transversal Clifford gates is both easier and less resource-intensive. In fact, finding codes accepting logical Clifford gates in a transversal manner is relatively straightforward. For example, any Calderbank–Shor–Steane (CSS) code constructed from a self-orthogonal binary cyclic code, or more generally doubly even self-orthogonal code, and its dual, admits transversal Clifford operators [12, 19, 36]. Nevertheless, those codes would lack the $T$-gate, implying that the quantum computations that can be realized would be classically simulable. A way to circumvent this problem is by means of magic state distillation and injection [7]. In this paradigm, magic state resources are injected into the “memory code” via a teleportation protocol resulting in the application of such operation on the code space in a fault-tolerant manner and the consumption of the magic resource [37]. The logical operation induced to the logical qubits will depend on the magic state resource injected, a common choice for the latter being |$\ket{T}$ magic states that result in the $T$ gate, which is the main non-Clifford gate discussed in this paper.
In order to apply the $T$ gate in a fault-tolerant manner, i.e., maintaining the error suppression level obtained by the code, magic states with a sufficient level of fidelity (of the order of the memory code) must be obtained. To do so, magic state distillation protocols such as the Bravyi–Haah protocols are used [4]. Such protocols make use of a certain amount of noisy magic states to distill a smaller number of magic states with higher fidelities. Magic state distillation protocols are closely connected to quantum error correction codes that admit the transversal implementation of a $T$-gate, i.e., such operation does not change the code space. These codes are commonly referred to as quantum codes that support transversal $T$.
Motivated by this application, numerous works in the literature have focused on constructing and characterizing stabilizer codes that support transversal $T$ gates; see, for instance, [1, 8, 9, 10, 32, 33]. In particular, [32] introduces CSS codes supporting transversal $T$ gates, known as CSS-T codes, which are the main object studied in this paper. The same authors demonstrate in [33] that CSS-T codes are optimal among all non-degenerate stabilizer codes supporting transversal $T$ gates.
However, a limitation of all currently known infinite families of CSS-T codes is that they exhibit either a vanishing rate or a vanishing relative distance. An intriguing open problem, proposed in [32], is to construct asymptotically good CSS-T codes, which simultaneously achieve a non-vanishing rate and relative distance.
Our contribution.
Motivated by the open problem mentioned above, we introduce a construction that produces a binary CSS-T code from any given binary CSS code. We employ this construction to prove the existence of asymptotically good binary CSS-T codes. Additionally, we show that our construction preserves the low-density parity check (LDPC) property, proving the existence of asymptotically good LDPC codes that are also CSS-T.
We also demonstrate that a generalization of our construction yields asymptotically good classes of binary codes that support transversal gates corresponding to any $Z$ rotation angle. Furthermore, we study the structure of the logical operators corresponding to the transversal $T$ gate, finer $Z$ rotations, and the $CCZ$ gate on our proposed CSS codes, which have the desired symmetry for these transversal gates.
Organization of the paper.
In Section 1, we set the notations and give definitions and known results on classical linear codes and quantum gates and codes. We also recall the definitions of CSS and CSS-T codes, which are the main focus of this paper. Section 2 presents our construction of CSS-T codes from CSS ones and their resulting parameters and properties. In particular, we prove here the existence of asymptotically good CSS-T codes. In Section 3, we discuss how to generalize our construction to support other transversal gates and examine their corresponding logical operators. Finally, in Section 4, we wrap up our work and present future research questions.
1 Preliminaries
In this section, we give a short introduction to classical and quantum error-correcting codes and their interplay. This allows us to fix the notations we will use for the rest of the paper. We refer the reader to [21] and to [17, 27, 35] for a detailed presentation of classical and quantum codes, respectively.
1.1 Classical linear codes
We work over the binary field $\mathbb{F}_{2}$. A (linear) code $C$ is an $\mathbb{F}_{2}$-linear subspace of $\mathbb{F}_{2}^{n}$, where $n$ is a positive integer. A vector $x\in C$ will be called a codeword. The length of the code $C$ is the dimension of the ambient space, namely $n$. The dimension of $C$ is its dimension as a linear subspace over $\mathbb{F}_{2}$. A linear code $C$ can be defined using a generator matrix, that is, a binary matrix with the row space equal to $C$. A linear code $D\subseteq C$ is called a subcode of $C$.
The support of a vector $x\in\mathbb{F}_{2}^{n}$ is the set ${\textrm{S}upp}(x):=\{i\in\{1,\dots,n\}\,|\,x_{i}\neq 0\}$. The (Hamming) weight of a codeword $x\in C$ is defined by ${\textrm{w}t}(x):=|{\textrm{S}upp}(x)|$. If ${\textrm{w}t}(x)\equiv 0\pmod{2}$, then $x$ is called an even (weight) vector. A linear code $C\subseteq\mathbb{F}_{2}^{n}$ is called an even code if it only contains even vectors. For a set of codewords $S$ we use the notation ${\textrm{w}t}(S)\coloneqq\{{\textrm{w}t}(x)\mid x\in S,x\neq\mathbf{0}\}$, where $\boldsymbol{0}$ denotes the all-zeros vector. Finally, the minimum distance $d(C)$ of a code $C$ is defined as
$d(C)\coloneqq\min\{{\textrm{w}t}(x)\mid x\in C,\,x\neq\boldsymbol{0}\}.$
We will refer to a code of length $n$, dimension $k$ and minimum distance $d$ as an $[n,k,d]$ code.
The (Euclidean) dual of $C\subseteq\mathbb{F}_{2}^{n}$ is the linear code
$C^{\perp}\coloneqq\{y\in\mathbb{F}_{2}^{n}\mid x\cdot y=0\mbox{ for all }x\in C\},$
where $x\cdot y$ denotes the Euclidean inner product of the vectors $x$ and $y$. If $C$ has dimension $k$, then $C^{\perp}$ has dimension $n-k$. A code $C$ is said to be self-orthogonal if $C\subseteq C^{\perp}$, and self-dual if $C=C^{\perp}$.
The component-wise multiplication of two length $n$ vectors $x$ and $y$, also known as the Schur product, is defined as
$(x_{1},x_{2},\ldots,x_{n})\star(y_{1},y_{2},\ldots,y_{n})\coloneqq(x_{1}y_{1},x_{2}y_{2},\ldots,x_{n}y_{n}).$
The Schur product of two codes $C,D\subseteq\mathbb{F}_{2}^{n}$ is then defined as
$C\star D\coloneqq{\textrm{S}pan}\{x\star y\mid x\in C,y\in D\}.$
The Schur product $C\star C$ is denoted $C^{\star 2}$. Note that for a binary linear code $C$, we have $x\star x=x$ for any $x\in C$, whence $C\subseteq C^{\star 2}$. We refer to [31] for an extensive survey on the Schur product of linear codes.
1.2 Quantum gates and quantum error correction
Throughout the paper, $\mathcal{H}$ is the complex Hilbert space $\mathbb{C}^{2}$ endowed with the inner product
$\langle u,v\rangle=\sum_{i=1}^{n}\overline{u_{i}}v_{i},$
where $v,u\in\mathcal{H}$ and $\overline{u_{i}}$ is the complex conjugate. A qubit of length one (or $1$-qubit) is an element of $\mathcal{H}$ of norm one. We show the set of one qubit Pauli operators by
$I=\begin{bmatrix}1&0\\ 0&1\end{bmatrix}$, $X=\begin{bmatrix}0&1\\ 1&0\end{bmatrix}$, $Y=\begin{bmatrix}0&-i\\ i&0\end{bmatrix}$, and $Z=\begin{bmatrix}1&0\\ 0&-1\end{bmatrix}$.
These operators form a basis for all $2\times 2$ matrices acting over $\mathcal{H}$. Hence an arbitrary one qubit error operator, i.e., $2\times 2$ unitary matrix over $\mathcal{H}$, can be represented as a linear combination of the mentioned matrices.
Let {|$\{\ket{0},\ket{1}\}$ be a basis for $\mathcal{H}$, where |$\ket{0}=\begin{bmatrix}1\\ 0\end{bmatrix}$ and |$\ket{1}=\begin{bmatrix}0\\ 1\end{bmatrix}$. An arbitrary state of a closed 1-qubit quantum system is defined by
|$\ket{\varphi}=\alpha\ket{0}+\beta\ket{1},$
[…]
Recall that the $Z$ rotation by the angle $\frac{2\pi}{2^{\ell}}$ for some $\ell\geq 0$ is given by the gate
$\varGamma^{(\ell)}=\begin{bmatrix}1&0\\ 0&e^{\frac{2\pi i}{2^{\ell}}}\end{bmatrix}.$
Then, the action of one qubit gate $\varGamma^{(\ell)}$ on the basis states is
$\begin{split}&\varGamma^{(\ell)}|0\rangle\longmapsto|0\rangle\ \text{and}\ \varGamma^{(\ell)}|1\rangle\longmapsto\mathrm{e}^{\frac{2\pi i}{2^{\ell}}}|1\rangle.\\ \end{split}$
If we fix $\ell=0,1,2,3$, then the $Z$ rotation $\varGamma^{(\ell)}$ is in correspondence to $I,Z,S$, and $T$, respectively.
For any $n\in\mathbb{N}$, we let $\mathcal{H}^{n}=\mathcal{H}\otimes\dots\otimes\mathcal{H}$ be the $2^{n}$-dimensional complex Hilbert space $\mathbb{C}^{2^{n}}$, where $\otimes$ denotes the usual tensor product of elements in $\mathbb{C}^{2}$. The inner product defined above for $\mathcal{H}$ generalizes to elements of $\mathcal{H}^{n}$. Qubits of length $n$ (or $n$-qubits) are elements of $\mathcal{H}^{n}$ of norm one. For the rest of this paper, we fix {|$\{\ket{a}\mid a\in\mathbb{F}_{2}^{n}\}$ to be a basis for $\mathcal{H}^{n}$.
1.3 From classical to quantum codes
We start by reviewing the construction of quantum CSS codes from a pair of classical codes. The following theorem summarizes the CSS construction of binary quantum codes, which is a commonly used technique in the literature.
Theorem 1 ([6]).
Let $C_{2},C_{1}\subseteq\mathbb{F}_{2}^{n}$ be two linear codes such that $C_{2}\subseteq C_{1}$. Then, there exists an $[\![n,k,d]\!]$ quantum code, called CSS and denoted by $(C_{1},C_{2})_{\mathrm{CSS}}$, where $k=\dim(C_{1})-\dim(C_{2})$ and $d=\min\{{\textrm{w}t}(C_{1}\setminus C_{2}),{\textrm{w}t}(C_{2}^{\perp}\setminus C_{1}^{\perp})\}$.
The quantum CSS code of Theorem 1 has a parity check matrix of the form
$\begin{bmatrix}H_{X}&0\\ 0&H_{Z}\end{bmatrix},$
where $H_{X}$ and $H_{Z}$ are generator matrices of $C_{2}$ and $C_{1}^{\perp}$, respectively.
Throughout the paper, we will use the notation $d_{X}$ and $d_{Z}$ for $\min\{{\textrm{w}t}(C_{1}\setminus C_{2})\}$ and $\min\{{\textrm{w}t}(C_{2}^{\perp}\setminus C_{1}^{\perp})\}$, respectively, and refer to them as the $X$-distance and the $Z$-distance, respectively. Note that since $d_{X}\geq\min\left\{d(C_{1})\right\}$ and $d_{Z}\geq\min\left\{d(C_{2}^{\bot})\right\}$, we always have $d\geq\min\left\{d(C_{1}),d(C_{2}^{\bot})\right\}$ for a quantum code $(C_{1},C_{2})_{\mathrm{CSS}}$. When equality holds in the latter inequality, we say that the quantum CSS code is non-degenerate. Otherwise, the code will be called degenerate.
Let $C_{2}\subseteq C_{1}$ be binary linear codes of length $n$ that define a binary quantum CSS code with parameters $[\![n,k,d]\!]$. Let $C_{1}=C_{2}\oplus{\textrm{S}pan}\{x_{1},x_{2},\ldots,x_{k}\}$, for some $x_{1},x_{2},\ldots,x_{k}\in C_{1}\setminus C_{2}$. Let $H$ be a $k\times n$ binary matrix having $x_{1},x_{2},\ldots,x_{k}$ as its rows. A common encoding approach for this binary quantum code is defined by ${\textrm{E}n}:\mathcal{H}^{k}\rightarrow\mathcal{H}^{n}$, where
${\textrm{E}n}\ket{u}=\frac{1}{\sqrt{|C_{2}|}}\sum_{x\in C_{2}}\ket{x+uH},$ (1)
where $u\in\mathbb{F}_{2}^{k}$. Note that the quantum state on the left-hand side of Equation (1), namely |$\ket{u}$, is a $k$-qubit state, while the right-hand side quantum states are $n$-qubit states. To distinguish them, from now on, we show the logical states using the notation |$\ket{u}_{L}$. Moreover, to simplify our computations, we discard the normalization factor. Each operation on such a quantum code is in the form of an $n$-qubit operation that preserves the code space, i.e., ${\textrm{E}n}(\mathcal{H}^{k})$.
In general, each error in an $n$-qubit operation can propagate within several of the code blocks. Therefore, a fault-tolerant approach is to investigate operations that can be decomposed into tensor product of $n$ single qubit gates, where each single error causes an error in one of the code blocks. Such $n$-qubit gates are called transversal gates. For instance, $T^{\otimes n}$ is an $n$-qubit transversal gate which will be called the transversal $T$ gate in this paper.
Recently, Rengaswamy et al. characterized all stabilizer codes in which transversal $T$ gate, and some other variants of it, preserve the code space [33]. In the literature, stabilizer codes with this property, specifically, those named triorthogonal, have been proposed for magic state distillation [4]. In [33], it also was shown that CSS codes are optimal among all non-degenerate stabilizer codes with this property. A restriction of their characterization to the CSS codes, commonly known as CSS-T codes, is stated below.
Definition 1 ([32]).
A CSS code $(C_{1},C_{2})_{\mathrm{CSS}}$ is called a CSS-T code if the followings hold:
  1. 1.
    the code $C_{2}$ is even, i.e., ${\textrm{w}t}(x)\equiv 0\pmod{2}$ for each $x\in C_{2}$;
  2. 2.
    for each $x\in C_{2}$, there exists a self-dual code in $C_{1}^{\perp}$ of dimension ${\textrm{w}t}(x)/2$ that is supported on $x$, i.e., there exists $C_{x}\subseteq C_{1}^{\perp}$ such that $|C_{x}|=2^{{\textrm{w}t}(x)/2}$, $C_{x}=C_{x}^{\perp_{x}}$, where $\perp_{x}$ denotes the dual taken in $\mathbb{F}_{2}^{{\textrm{w}t}(x)}$, and for each $z\in C_{x}$ we have ${\textrm{S}upp}(z)\subseteq{\textrm{S}upp}(x)$.
A reformulation of the above conditions in terms of the Schur product of the ingredient classical codes was proved in [9] and we recalled below.
Theorem 2 ([9, Theorem 2.3]).
A pair of binary linear codes $C_{1}$ and $C_{2}$ generates a CSS-T code if and only if $C_{1}^{\perp}+C_{1}^{\star 2}\subseteq C_{2}^{\perp}$.
It is worth noting that if we consider a pair of linear codes $C_{2}\subseteq C_{1}$ giving a CSS code, we obtain for free the inclusion $C_{1}^{\perp}\subseteq C_{2}^{\perp}$. Therefore, each given CSS code $(C_{1},C_{2})_{\mathrm{CSS}}$, is a CSS-T code if and only if $C_{1}^{\star 2}\subseteq C_{2}^{\perp}$.
Some known constructions and several examples of CSS-T codes are provided in [1, 8, 9, 10, 32, 33]. All currently known families of CSS-T codes have either a vanishing rate or a vanishing relative distance. An interesting open problem, which will be answered in the next section, is to construct an infinite family of asymptotically good CSS-T codes with non-vanishing rate and relative distance [32].
2 Constructing CSS-T codes from CSS codes
In this section, we present a method to construct a CSS-T code from any given CSS code. This allows us to prove in Theorem 4 and Corollary 1 the existence of families of asymptotically good CSS-T codes. In particular, our construction preserves the LDPC property of the original CSS code, enabling us to show in Theorem 5 and Corollary 2 the existence of asymptotically good quantum LDPC codes that are CSS-T.
Our approach is based on a technique that doubles the ingredient pair of classical codes. This entire section builds up to the following result.
Theorem 3.
For any given $[\![n,k,d]\!]$ binary CSS code $(C_{1},C_{2})_{\mathrm{CSS}}$, one can construct a $[\![2n,k,\geq d]\!]$ binary CSS-T code.
In the remainder of this section, we provide more details about our construction and prove the above mentioned result.
Definition 2.
Let $C\subseteq\mathbb{F}_{2}^{n}$ be a linear code. Let $\phi:C\to\mathbb{F}_{2}^{n}$ be a map. We define
$C^{N_{\phi}}\vcentcolon=\left\{(x,\phi(x))\mid x\in C\right\}.$
In what follows, we focus on using a CSS code $(C_{1},C_{2})_{\mathrm{CSS}}$ and an appropriate map $\phi$ such that $\left(C_{1}^{N_{\phi}},C_{2}^{N_{\phi}}\right)$ forms a CSS-T code. In the following proposition, we give necessary and sufficient conditions on the map $\phi$ so that $\left(C_{1}^{N_{\phi}},C_{2}^{N_{\phi}}\right)$ forms a CSS-T code.
Proposition 1.
Let $(C_{1},C_{2})_{\mathrm{CSS}}$ be a CSS code of length $n$. Then $\left(C_{1}^{N_{\phi}},C_{2}^{N_{\phi}}\right)$ forms a CSS-T code if and only if $\phi:C_{1}\rightarrow\mathbb{F}_{2}^{n}$ is a linear map such that for each $x,y\in C_{1}$ and $z\in C_{2}$ we have
${\textrm{w}t}(x\star y\star z)+{\textrm{w}t}(\phi(x)\star\phi(y)\star\phi(z))\equiv 0\mod{2}.$
Proof.
By definition, $C_{2}^{N_{\phi}}\subseteq C_{1}^{N_{\phi}}$. Hence, $\left(C_{1}^{N_{\phi}},C_{2}^{N_{\phi}}\right)$ is a CSS code if and only if $C_{1}^{N_{\phi}}$ and $C_{2}^{N_{\phi}}$ are linear codes, that is, if and only if for any $x,y\in C_{1}$ we have $(x,\phi(x))+(y,\phi(y))=(x+y,\phi(x)+\phi(y))=(x+y,\phi(x+y))$, where the second equality follows by the definition of $C_{1}^{N_{\phi}}$. Thus, $\phi$ must be linear on $C_{1}$. Now, $\left(C_{1}^{N_{\phi}},C_{2}^{N_{\phi}}\right)_{\mathrm{CSS}}$ is CSS-T if and only if $\left({C_{1}^{N_{\phi}}}\right)^{\star 2}\subseteq\left({C_{2}^{N_{\phi}}}\right)^{\perp}$ by Theorem 2. So $\left(C_{1}^{N_{\phi}},C_{2}^{N_{\phi}}\right)_{\mathrm{CSS}}$ is CSS-T if and only if for each $x,y\in C_{1}$ and each $z\in C_{2}$ we have
$((x,\phi(x))\star(y,\phi(y)))\cdot(z,\phi(z))$ $=(x\star y)\cdot z+(\phi(x)\star\phi(y))\cdot\phi(z)$
$=\sum_{i=1}^{n}x_{i}y_{i}z_{i}+\sum_{i=1}^{n}\phi(x)_{i}\phi(y)_{i}\phi(z)_{i}=0.$
The last equality holds if and only if ${\textrm{w}t}(x\star y\star z)+{\textrm{w}t}(\phi(x)\star\phi(y)\star\phi(z))\equiv 0\pmod{2}$. This completes the proof. ∎
Proposition 2.
Let $\phi$ be a map satisfying the conditions of Proposition 1. If $(C_{1},C_{2})_{\mathrm{CSS}}$ is an $[\![n,k]\!]$ CSS code, then $\left(C_{1}^{N_{\phi}},C_{2}^{N_{\phi}}\right)$ is a $[\![2n,k]\!]$ CSS-T code.
Proof.
This follows immediately by the properties of $\phi$ and the definition of $C^{N_{\phi}}$. ∎
Any map $\phi$ satisfying the conditions of Proposition 1 allows constructing a CSS-T code from a CSS one. Let us point out that permutations of the coordinates, including the identity, satisfy those properties. Furthermore, using different permutations as the $\phi$ map will result in CSS codes with the same parameters, as shown in the following proposition.
Proposition 3.
Let $C_{2}\subseteq C_{1}\subseteq\mathbb{F}_{2}^{n}$ be a pair of binary codes, and let $I$ and $\phi$ be the identity map and a permutation map on $\mathbb{F}_{2}^{n}$, respectively. Then, the CSS codes $\left(C_{1}^{N_{I}},C_{2}^{N_{I}}\right)_{\mathrm{CSS}}$ and $\left(C_{1}^{N_{\phi}},C_{2}^{N_{\phi}}\right)_{\mathrm{CSS}}$ have the same parameters.
Proof.
Clearly, the two CSS codes have the same length and the same dimension. So, we only need to deal with the minimum distance. Let $\Phi=(I,\phi)$ be the permutation of $\mathbb{F}_{2}^{2n}$ that leaves the first $n$ coordinates unchanged and applies $\phi$ to the second $n$ coordinates. Then $\Phi\left(C_{1}^{N_{I}}\right)=C_{1}^{N_{\phi}}$ and $\Phi\left(C_{2}^{N_{I}}\right)=C_{2}^{N_{\phi}}$. Therefore, $\Phi\left(C_{1}^{N_{I}}\setminus C_{2}^{N_{I}}\right)=C_{1}^{N_{\phi}}\setminus C_{2}^{N_{\phi}}$. Hence, both CSS codes have the same $X$-distance.
Since $\Phi$ is a permutation, we have $\Phi\left(\left({C_{i}^{N_{I}}}\right)^{\bot}\right)=\left(\Phi\left(C_{i}^{N_{I}}\right)\right)^{\bot}$ for $i\in\{1,2\}$. Thus $\Phi\left(\left({C_{i}^{N_{I}}}\right)^{\bot}\right)=\left(\Phi\left(C_{i}^{N_{I}}\right)\right)^{\bot}=\left(C_{i}^{N_{\phi}}\right)^{\bot}$ and $\Phi\left({\left(C_{2}^{N_{I}}\right)}^{\bot}\setminus{\left(C_{1}^{N_{I}}\right)}^{\bot}\right)={\left(C_{2}^{N_{\phi}}\right)}^{\bot}\setminus{\left(C_{1}^{N_{\phi}}\right)}^{\bot}$ for $i\in\{1,2\}$. Hence, both CSS codes have the same $Z$-distance. ∎
In the rest of the section, we fix $\phi$ to be the identity permutation $I$, and consider the code
$C^{N_{I}}\vcentcolon=\{(x,x)\mid x\in C\},$ (2)
which we will denote simply by $C^{N}$.
Remark 1.
Other choices of $\phi$ are possible and may lead to CSS-T codes with different parameters. As an example, consider the map $\phi:x\mapsto x+a$ for some $a\in C_{1}\cap(C_{1}^{\star 2})^{\perp}$. Finding maps $\phi$ giving CSS-T codes with better parameters than the choice $\phi=I$ is an interesting open question. We return to this discussion in the conclusion.
By Proposition 2 we know that $\left(C_{1}^{N},C_{2}^{N}\right)_{\mathrm{CSS-T}}$ has length $2n$ and dimension $k=\dim(C_{1})-\dim(C_{2})$. Hence, in what follows, we focus on studying the minimum distance of this code. To this end, we set $d^{N}=\min\left\{d_{X}^{N},d_{Z}^{N}\right\}$ for the CSS-T code $\left(C_{1}^{N},C_{2}^{N}\right)_{\mathrm{CSS-T}}$, where $d_{X}^{N}=\min\left\{{\textrm{w}t}\left(C_{1}^{N}\setminus C_{2}^{N}\right)\right\}$ and $d_{Z}^{N}=\min\left\{{\textrm{w}t}\left((C_{2}^{N})^{\perp}\setminus(C_{1}^{N})^{\perp}\right)\right\}$.
Lemma 1.
The CSS-T code $\left(C_{1}^{N},C_{2}^{N}\right)_{\mathrm{CSS}}$ satisfies $d_{X}^{N}=2d_{X}$.
Proof.
Let $x^{N}=(x,x)$ be the minimum weight codeword of $C_{1}^{N}\setminus C_{2}^{N}$, where $x\in C_{1}$. Note that $x\in C_{1}\setminus C_{2}$ as otherwise $x^{N}\in C_{2}^{N}$. The fact that $x^{N}$ has the minimum weight implies that $x$ has to have weight $d_{X}$. Therefore, ${\textrm{w}t}(x^{N})=2d_{X}$ and we have $d_{X}^{N}=2d_{X}$. ∎
To calculate $d_{Z}^{N}$, we need the following technical lemma.
Lemma 2.
The followings hold.
  1. 1.
    Let $x=(x_{1},x_{2})\in\left(C_{2}^{N}\right)^{\perp},$ for some $x_{i}\in\mathbb{F}_{2}^{n}$. Then $x^{\prime}=(x_{1}+x_{2},0)\in\left(C_{2}^{N}\right)^{\perp}$. In particular, $x_{1}+x_{2}\in C_{2}^{\bot}$ and ${\textrm{w}t}(x^{\prime})\leq{\textrm{w}t}(x)$.
  2. 2.
    For each $x=(a,a)\in\left(C_{2}^{N}\right)^{\perp}$, we have $x\in\left(C_{1}^{N}\right)^{\perp}$.
  3. 3.
    Let $x_{1},x_{2}\in\mathbb{F}_{2}^{n}$. Then $x_{1}+x_{2}\in C_{1}^{\perp}$ if and only if $x^{\prime}=(x_{1}+x_{2},0)\in\left(C_{1}^{N}\right)^{\perp}$.
Proof.
(1) For each $y=(a,a)\in C_{2}^{N}$ we have
$x\cdot y=x_{1}\cdot a+x_{2}\cdot a=(x_{1}+x_{2})\cdot a=x^{\prime}\cdot y=0.$
The fact that $x_{1}+x_{2}\in C_{2}^{\bot}$ and ${\textrm{w}t}(x^{\prime})\leq{\textrm{w}t}(x)$ follows immediately.
(2) Let $y=(b,b)\in C_{1}^{N}$. Then $x\cdot y=a\cdot b+a\cdot b=0$. Hence $x\in\left(C_{1}^{N}\right)^{\perp}$.
(3) The statement is proved by the following sequence of equivalences:
$x_{1}+x_{2}\in C_{1}^{\perp}$ $\iff\forall a\in C_{1},(x_{1}+x_{2})\cdot a=0\iff\forall y=(a,a)\in C_{1}^{N},x^{\prime}\cdot y=0$
$\iff x^{\prime}=(x_{1}+x_{2},0)\in\left(C_{1}^{N}\right)^{\perp}.$
Next, we use the above lemma to compute $d_{Z}^{N}$.
Proposition 4.
The CSS-T code $\left(C_{1}^{N},C_{2}^{N}\right)_{\mathrm{CSS}}$ satisfies $d_{Z}^{N}=d_{Z}$.
Proof.
Let $x=(x_{1},x_{2})\in\left(C_{2}^{N}\right)^{\perp}\setminus\left(C_{1}^{N}\right)^{\perp}$ for some $x_{1},x_{2}\in\mathbb{F}_{2}^{n}$ be a minimum weight vector. Since $d_{Z}^{N}={\textrm{w}t}\left(\left(C_{2}^{N}\right)^{\perp}\setminus\left(C_{1}^{N}\right)^{\perp}\right)$, Lemma 2(2) implies that $x_{1}\neq x_{2}$. Next, Lemma 2(2) implies that $x^{\prime}=(x_{1}+x_{2},0)\in\left(C_{2}^{N}\right)^{\perp}$. Moreover, the fact that $x\notin\left(C_{1}^{N}\right)^{\perp}$ implies the existence of $c\neq 0\in C_{1}$ such that $x\cdot(c,c)=x_{1}\cdot c+x_{2}\cdot c\neq 0$. Therefore, $x^{\prime}\cdot(c,c)\neq 0$ which implies $x^{\prime}\notin\left(C_{1}^{N}\right)^{\perp}$. Now, Lemma 2(3) implies that $x_{1}+x_{2}\in\left(C_{2}^{\perp}\setminus C_{1}^{\perp}\right)$. This shows that
$d_{Z}^{N}={\textrm{w}t}(x)\geq{\textrm{w}t}(x^{\prime})={\textrm{w}t}(x_{1}+x_{2})\geq d_{Z}.$ (3)
For the other inequality, let $x$ be a vector of minimum weight in $C_{2}^{\perp}\setminus C_{1}^{\perp}$. Then, observe that $(x,0)\in\left(C_{2}^{N}\right)^{\perp}\setminus\left(C_{1}^{N}\right)^{\perp}$ as for each $(y,y)\in C_{2}^{N}$, where $y\in C_{2}$, we have $(x,0)\cdot(y,y)=x\cdot y=0.$ The fact that $(x,0)\not\in(C_{1}^{N})^{\perp}$ follows from Lemma 2(3). Thus
${\textrm{w}t}((x,0))=d_{Z}\geq d_{Z}^{N}.$ (4)
Finally, Equations (3) and (4) complete the proof. ∎
Finally, we are ready to prove Theorem 3.
Proof of Theorem 3.
Let $(C_{1},C_{2})_{\mathrm{CSS}}$ have parameters $[\![n,k,d]\!]$. The combination of Propositions 2 and 4 and Lemma 1 shows that the CSS-T family $(C_{1}^{N},C_{2}^{N})$ has parameters $[\![2n,k,\geq d]\!]$. This concludes the proof. ∎
Remark 2.
Let $(C_{1},C_{2})_{\mathrm{CSS}}$ be a CSS code with the $X$ and $Z$ distances $d_{X}$ and $d_{Z}$, respectively.
  1. 1.
    When $d_{X}<d_{Z}$ the CSS-T code of Theorem 3 has a strictly larger minimum distance than the initial CSS code. On the other hand, if $d_{X}>d_{Z}$ the CSS-T code has the same minimum distance as the initial CSS code.
  2. 2.
    When $d_{X}\neq d_{Z}$, one can apply the result of Theorem 3 to construct CSS-T code with double length, improved minimum distance and the same dimension of the ingredient CSS code. In particular, if $d_{X}<d_{Z}$ in a CSS code $(C_{1},C_{2})_{CSS}$, then applying Theorem 3 directly to this code gives the desired CSS-T code. In the case when $d_{X}>d_{Z}$, then one can consider $C_{1}^{\prime}=C_{2}^{\perp}$ and $C_{2}^{\prime}=C_{1}^{\perp}$, after exchanging the $X$ and $Z$ stabilizers in the CSS code. Then, applying Theorem 3 to $C_{2}^{\prime}\subseteq C_{1}^{\prime}$ gives the desired result.
  3. 3.
    If $d_{X},d_{Z}>2$, then the CSS-T code of Theorem 3 is a degenerate quantum code. In particular, it contains all weight two codewords in the form $(a,a)$, where $a\in\mathbb{F}_{2}^{n}$ is a weight one vector.
Table 1 provides a list of examples of binary CSS-T codes that we found after applying the construction of Theorem 3 to binary linear cyclic codes.
CSS-T Parameters
$[\![14,3,3]\!]$ $[\![18,2,3]\!]$ $[\![30,1,6]\!]$ $[\![30,2,6]\!]$
$[\![30,8,4]\!]$ $[\![30,10,3]\!]$ $[\![42,1,6]\!]^{\ast}$ $[\![42,6,6]\!]$
$[\![42,7,5]\!]$ $[\![42,12,4]\!]$ $[\![46,1,7]\!]$ $[\![54,1,6]\!]^{\ast}$
$[\![62,1,11]\!]$ $[\![62,15,6]\!]$ $[\![66,2,10]\!]^{\ast}$
Table 1: Small length binary CSS-T codes from cyclic codes. All the CSS-T codes are $Z$-degenerate. The codes shown by $\ast$ are also $X$-degenerate.
2.1 Asymptotically good families of CSS-T codes
Recall that the rate and relative distance of a binary quantum code with parameters $[\![n,k,d]\!]$ are defined as $R=\frac{k}{n}$ and $\delta=\frac{d}{n}$, respectively. In the literature there exist several infinite families of quantum CSS codes with non-vanishing asymptotic rate and relative distance, commonly known as asymptotically good quantum codes, see for example [2, Theorem 1], [11, Theorem 1.2], and [28, Theorem 2]. The existence of such families combined with our Theorem 3 allow us to construct asymptotically good CSS-T codes.
Theorem 4.
From any given asymptotically good family of binary CSS codes it is possible to derive an asymptotically good family of binary CSS-T codes.
Proof.
Let $\{(C_{1},C_{2})_{i}\}$ be a family of asymptotically good CSS codes, where $i>0$ is an integer, with rate and relative distance $\frac{k_{i}}{n_{i}}$ and $\frac{d_{i}}{n_{i}}$, respectively, for each $i$. Let the asymptotic rate and relative distance be $R=\displaystyle\lim_{i\to\infty}\frac{k_{i}}{n_{i}}>0$ and $\delta=\displaystyle\lim_{i\to\infty}\frac{d_{i}}{n_{i}}>0.$ Then applying the result of Theorem 3 gives the infinite family $\left\{\left({C_{1}}^{N},{C_{2}}^{N}\right)_{i}\right\}$ for each $i>0$ with asymptotic rate and relative distance $R^{\prime}=\frac{R}{2}>0$ and $\delta^{\prime}=\frac{\delta}{2}>0$, respectively. ∎
Corollary 1.
Infinite families of quantum CSS-T codes with non-vanishing asymptotic rate and relative distance exist.
Proof.
Examples of infinite asymptotically good families of binary CSS codes are presented, for instance, in [2, Theorem 1], [11, Theorem 1.2], and [28, Theorem 2]. The result follows from Theorem 4. ∎
2.2 Asymptotically good families of LDPC CSS-T codes
In this subsection, we show that the CSS-T construction of Theorem 3 preserves the Low-Density Parity Check (LDPC) property. An immediate application of this is the existence of asymptotically good quantum LDPC CSS-T codes. Quantum codes with sparse parity check matrices are especially interesting in the context of quantum computing. The LDPC property implies that the computation of each parity constraint imposed by the code involves a small number of qubits. As a result, the qubits of the code need to interact by means of a small number of two-qubit gates. Due to the fact that those gates used for syndrome extraction are also noisy, having a sparse amount of parity constraints is beneficial for practical implementations of quantum error correction codes [3, 5]. Furthermore, sparse parity check matrices are also important for decoding purposes [13, 34]. Thus, we discuss how our construction can be used to yield asymptotically good quantum LDPC codes. Before stating this result, we need some basic definitions and elementary results.
Let $C_{2}\subseteq C_{1}\subseteq\mathbb{F}_{2}^{n}$ be binary linear codes. Recall that the corresponding quantum CSS code has a parity check matrix
$H=\begin{bmatrix}H_{X}&0\\ 0&H_{Z}\end{bmatrix},$
where $H_{X}$ and $H_{Z}$ are generator matrices of $C_{2}$ and $C_{1}^{\perp}$, respectively. A quantum code with a sparse parity-check matrix $H$ is called a quantum LDPC code. Recall also that the following definition was given in Equation (2):
$C_{i}^{N_{I}}=\left\{(x,x):x\in C_{i}\right\}$
for each $i\in\{1,2\}$. In the following, we will use the shorter notation $C_{i}^{N}$ for $C_{i}^{N_{I}}$.
The next theorem shows that applying the result of Theorem 3 to a CSS code preserves its sparsity by sending its parity check matrix to another sparse parity check matrix.
Theorem 5.
Let $(C_{1},C_{2})_{\mathrm{CSS}}$ be a CSS code, and let
$H=\begin{bmatrix}H_{X}&0\\ 0&H_{Z}\end{bmatrix},$
where $H_{X},H_{Z}$ generate $C_{2}$ and $C_{1}^{\perp}$, respectively. Then, the CSS-T code $\left(C_{1}^{N},C_{2}^{N}\right)$ has parity check matrix
$H^{N}=\begin{bmatrix}H_{X}^{N}&0\\ 0&H_{Z}^{N}\end{bmatrix},$
where
$H^{N}_{X}=\begin{bmatrix}H_{X}&H_{X}\end{bmatrix}$ and $H^{N}_{Z}=\begin{bmatrix}H_{Z}&0\\ I&I\end{bmatrix}.$
In particular, if $r_{X}$ and $r_{Z}$ are the maximum row weights of $H_{X}$ and $H_{Z}$, respectively, then the maximum row weights of $H_{X}^{N}$ and $H_{Z}^{N}$ are $2r_{X}$ and $\max\{r_{Z},2\}$.
Proof.
The definition of $C_{2}^{N}$ justifies that $H_{X}^{N}$ is a generator matrix for $C_{2}^{N}$. Moreover, one can easily check that the linear span of the rows of $H_{Z}^{N}$ is a subspace of $\left(C_{1}^{N}\right)^{\bot}$. Next, we show that the row space of $H_{Z}^{N}$ is in fact $(C_{1}^{N})^{\bot}$. Let $(x,y)\in\left(C_{1}^{N}\right)^{\perp}$ for some $x,y\in\mathbb{F}_{2}^{n}$. Let $e_{i}^{N}\vcentcolon=(e_{i},e_{i})$, where $e_{i}\in\mathbb{F}_{2}^{n}$ is the standard basis vector whose $j$-th entry is
$({e_{i}})_{j}=\begin{cases}1&\text{ if }i=j,\\ 0&\text{ if }i\neq j.\end{cases}$
So there exists $\alpha_{i}\in\mathbb{F}_{2}$ for each $1\leq i\leq n$ such that
$(x,y)+\sum_{i=1}^{n}\alpha_{i}e_{i}^{N}=(x,y)+(y,y)=(x+y,0)\in\left(C_{1}^{N}\right)^{\bot}.$
Now Lemma 2(3) implies that $x+y\in C_{1}^{\bot}$ and consequently $(x+y,0)$ is in the row space of $\begin{bmatrix}H_{Z}&0\end{bmatrix}$. Since $(y,y)$ belongs to the row space of $\begin{bmatrix}I&I\end{bmatrix}$, we conclude that $(x,y)$ is in the row space of $H_{Z}^{N}$.
The structure of $H_{X}^{N}$ shows that its maximum row weight is $2r_{X}$. Moreover, the last $n$ rows of $H_{Z}^{n}$ all have weight $2$. The maximum weight of the remaining rows of $H_{Z}^{N}$ is $r_{Z}$. Thus the maximum row weight of $H_{Z}^{N}$ is $\max\{r_{Z},2\}$. ∎
Recall that there exist several families of good quantum LDPC codes in the literature, for instance, see [14, 16, 23, 24, 29, 30]. An immediate consequence of the above theorem is that applying the construction of Theorem 3 to an LDPC CSS code preserves its sparsity. We use this property to entail the next result.
Corollary 2.
A family of asymptotically good quantum LDPC code that is CSS-T exists.
Proof.
The proof follows by applying the quantum CSS-T construction of Theorem 3 to each asymptotically good quantum LDPC family presented above. Theorem 5 ensures that the resulting CSS-T codes are LDPC. ∎
3 Logical operators of different transversal gates
In this section, we first develop the idea used in the previous chapter to produce asymptotically good CSS codes for any given $Z$ rotation. Secondly, we show that for any $(C_{1},C_{2})_{\mathrm{CSS-T}}$ with an even code $C_{1}$, the logical operator corresponding to the transversal $T$ has multiplicative order four. Therefore, the transversal $T$ gate never realizes transversal logical $T$ gate on such CSS codes. Finally, we prove that the CSS-T codes from the construction of Theorem 3 also support transversal $CCZ$ gate. In this case, the corresponding logical operator is the identity.
3.1 Good CSS codes supporting transversal Z rotations
As we already stated in Corollary 1, one can construct a family of asymptotically good CSS codes that supports transversal $T$ gate. Obviously, the same family supports other larger $Z$ rotations, namely transversal $Z$ or $S$ gate. In this subsection, we discuss the construction of another asymptotically good CSS family that supports transversal application of other smaller $Z$ rotation gates.
First, recall that the $Z$ rotation gate by the angle $\frac{2\pi}{2^{\ell}}$ for some $\ell\geq 0$ is
$\varGamma^{(\ell)}=\begin{bmatrix}1&0\\ 0&e^{\frac{2\pi i}{2^{\ell}}}\end{bmatrix}.$
Next, we show that applying the construction of Theorem 3 repeatedly preserves finer $Z$ rotations transversally.
Theorem 6.
Let $\ell>3$ be an integer. Applying transversal $\varGamma^{(\ell)}$ gate to binary quantum CSS code $\left(C_{1}^{\ell},C_{2}^{\ell}\right)$, obtained from applying Theorem 3 $\ell$ times to $C_{2}\subseteq C_{1}\subseteq\mathbb{F}_{2}^{n}$, results in the logical identity operator.
Proof.
Let $u\in\mathbb{F}_{2}^{k}$, where $k$ is the number of logical qubits in the quantum code, and $H$ be a binary matrix having a basis of $C_{3}^{\ell}$ in its rows, where $C_{1}^{\ell}=C_{2}^{\ell}\oplus C_{3}^{\ell}$. Note that the number of physical qubits is $2^{\ell}n$. Next, we show that the following diagram commutes
|${\ket{u}_{L}}$${\displaystyle\sum_{v\in C_{2}^{l}}\ket{v+uH}}$|${\ket{u}_{L}}$${\displaystyle\sum_{v\in C_{2}^{l}}\varGamma^{(\ell)^{\otimes{2^{l}n}}}\ket{v+uH}}$$\scriptstyle{{\textrm{E}n}}$$\scriptstyle{I}$$\scriptstyle{\varGamma^{(\ell)^{\otimes{2^{\ell}n}}}}$$\scriptstyle{{\textrm{E}n}}$ (5)
The action of applying $\varGamma^{(\ell)}$ to all $2^{\ell}n$ physical qubits can be described as:
$\varGamma^{(\ell)^{\otimes{2^{\ell}n}}}\left(\sum_{v\in C_{2}^{\ell}}|v+uH\rangle\right)=\sum_{v\in C_{2}^{\ell}}\varGamma^{(\ell)^{\otimes{2^{\ell}n}}}|v+uH\rangle.$ (6)
Recall that the term $v+uH\in C_{1}^{\ell}$ for each selection of $u$ and $v$. Now, applying the operator $\varGamma^{(\ell)^{\otimes{2^{\ell}n}}}$ to each term in Equation (6) gives
$\sum_{v\in C_{2}^{\ell}}\varGamma^{(\ell)^{\otimes{2^{\ell}n}}}|v+uH\rangle=\sum_{v\in C_{2}^{\ell}}\left(\mathrm{e}^{\frac{2\pi i}{2^{\ell}}}\right)^{{\textrm{w}t}{(v+uH)}}|v+uH\rangle,$
where ${\textrm{w}t}{(v+uH)}$ is the weight of corresponding vector in $C_{1}^{\ell}$. Moreover, since $C_{1}^{\ell}$ is obtained from applying the procedure of Theorem 3 $\ell$ times, its codewords are in the form
$C_{1}^{\ell}=\{(x,x,\ldots,x):x\in C_{1}\},$
where the number of repetitions is $2^{\ell}$ times. Hence the weight of each codeword of $C_{1}^{\ell}$ is divisible by $2^{\ell}$. Thus
$\sum_{v\in C_{2}^{\ell}}\varGamma^{(\ell)^{\otimes{2^{\ell}n}}}|v+uH\rangle=\sum_{v\in C_{2}^{\ell}}\left(\mathrm{e}^{\frac{2\pi i}{2^{l}}}\right)^{{\textrm{w}t}{(v+uH)}}|v+uH\rangle=\sum_{v\in C_{2}^{\ell}}|v+uH\rangle=|u\rangle_{L}.$
This shows that the diagram of Equation (5) commutes, and completes the proof. ∎
Note that one may get the same result, with a different logical operator, after applying Theorem 3 to $C_{2}\subseteq C_{1}$ less than $\ell$ times. However, using $\ell$ repetitions keeps the proof more intuitive and simpler. An application of Theorem 6 is the following interesting result. The proof is similar to that of Corollary 1. So we skip it in here.
Corollary 3.
Let $\ell>0$ be a positive integer. Then a family of asymptotically good CSS code exists that supports transversal $\varGamma^{(\ell)}$ gate.
3.2 Logical operator of CSS-T codes with even weights
Recall that applying the $T$ gate transversally to all the physical qubits of a CSS-T code realizes a logical gate. In this subsection, we investigate the structure of this logical operator when the $C_{2}\subseteq C_{1}$ both are even and form a CSS-T code. We show that the multiplicative order of such a logical operator is always a divisor of four. In particular, such CSS-T codes never realize a transversal logical $T$ gate as it has the multiplicative order of eight.
Theorem 7.
Let $(C_{1},C_{2})$ be a CSS-T code with length $n$ such that ${\textrm{w}t}(x)\equiv 0\pmod{2}$ for each $x\in C_{1}$. Then, the logical operator corresponding to $T^{\otimes n}$ has order a factor of four.
Proof.
Let $H$ be a binary matrix having a basis of $C_{3}$ in its rows, where $C_{1}=C_{2}\oplus C_{3}$. The CSS code of $(C_{1},C_{2})$ preserves the transversal $T$, implying that the following diagram commutes:
|${\ket{u}_{L}}$${\displaystyle\sum_{v\in C_{2}}\ket{v+uH}}$|${\ket{u^{\prime}}_{L}}$${\displaystyle\sum_{v\in C_{2}}\ket{v+u^{\prime}H}=\displaystyle\sum_{v\in C_{2}}T^{\otimes n}\ket{v+uH},}$$\scriptstyle{{\textrm{E}n}}$$\scriptstyle{OP}$$\scriptstyle{T^{\otimes n}}$$\scriptstyle{{\textrm{E}n}}$ (7)
where $OP$ is the corresponding logical operator. Applying the operator $OP$ four times implies the next diagram
[…]
Moreover,
${T^{4}}^{\otimes n}\ket{v+uH}=\mathrm{e}^{(\frac{8\pi i}{8}){\textrm{w}t}(v+uH)}\ket{v+uH}=\ket{v+uH}$
as ${\textrm{w}t}(v+uH)$ is even. So the bottom-left part of Equation (8) should be the same as the top-left part (i.e., |$\ket{u}_{L}=\ket{u^{\prime\prime}}_{L}$), which implies that $OP^{4}$ is the logical identity. ∎
Recall that $C_{2}^{N}\subseteq C_{1}^{N}$ is the pair of classical codes used in Theorem 3. Their definition shows that $C_{1}^{N}$ also has only even weight codewords. Therefore, the logical operator corresponding to the transversal $T$, when applied to the quantum code of Theorem 3, has order a factor of four. All the CSS-T codes obtained from the extended construction of classical codes and the CSS-T Reed–Muller (RM) (but not punctured RM) codes satisfy the conditions of Theorem 7.
3.3 Application of transversal CCZ gate
Let $u,v,w\in\mathbb{F}_{2}^{n}$. Recall that the transversal $CCZ$ gate is defined as follows:
$CCZ^{\otimes n}|u\rangle|v\rangle|w\rangle\longmapsto(-1)^{\displaystyle\sum_{i=1}^{n}(u_{i}v_{i}w_{i})}|u\rangle|v\rangle|w\rangle.$
In this subsection, we show that the quantum CSS code obtained from Theorem 3 also support transversal CCZ gate. First, recall that for each $C_{2}\subseteq C_{1}$, we have set
$C_{i}^{N}\vcentcolon=\{(z,z):z\in C_{i}\}$
for each $i\in\{1,2\}$. We also fix $H$ to be a binary matrix having a basis of $C_{3}^{N}$ in its rows, where $C_{1}^{N}=C_{2}^{N}\oplus C_{3}^{N}$.
Theorem 8.
Let $C_{2}\subseteq C_{1}\subseteq\mathbb{F}_{2}^{n}$ be two binary codes. The quantum CSS-T code $\left(C_{1}^{N},C_{2}^{N}\right)_{\mathrm{CSS-T}}$ has the following property: applying the transversal $CCZ$ gate to all $2n$ physical qubits realizes the logical identity gate.
Proof.
[…]
$CCZ^{\otimes 2n}\left(\sum_{v\in C_{2}^{N}}|v+uH\rangle\right)\left(\sum_{v^{\prime}\in C_{2}^{N}}|v^{\prime}+xH\rangle\right)\left(\sum_{v^{\prime\prime}\in C_{2}^{N}}|v^{\prime\prime}+wH\rangle\right).$ (9)
Expanding this gives:
$CCZ^{\otimes 2n}\left(\sum_{v,v^{\prime},v^{\prime\prime}\in C_{2}^{N}}|v+uH\rangle|v^{\prime}+xH\rangle|v^{\prime\prime}+wH\rangle\right).$
Now, applying $CCZ^{\otimes 2n}$ to each term in the summation gives:
$\begin{split}&\sum_{v,v^{\prime},v^{\prime\prime}\in C_{2}^{N}}CCZ^{\otimes 2n}|v+uH\rangle|v^{\prime}+xH\rangle|v^{\prime\prime}+wH\rangle\\ &=\sum_{v,v^{\prime},v^{\prime\prime}\in C_{2}^{N}}CCZ^{\otimes 2n}|a\rangle|b\rangle|c\rangle=\sum_{v,v^{\prime},v^{\prime\prime}\in C_{2}^{N}}(-1)^{\displaystyle\sum_{i=1}^{2n}a_{i}b_{i}c_{i}}|a\rangle|b\rangle|c\rangle,\end{split}$
where $a=v+uH=(\alpha,\alpha)$, $b=v^{\prime}+xH=(\beta,\beta)$, and $c=v^{\prime\prime}+wH=(\gamma,\gamma)$ are codewords of $C_{1}^{N}$ for some $\alpha,\beta,\gamma\in\mathbb{F}_{2}^{n}$. Thus
$\sum_{i=1}^{2n}a_{i}b_{i}c_{i}=2\sum_{i=1}^{n}\alpha_{i}\beta_{i}\gamma_{i}\equiv 0\pmod{2}.$
This implies that $(-1)^{\displaystyle\sum_{i=1}^{2n}a_{i}b_{i}c_{i}}=1$. Thus, the result of (9) is
$\left(\sum_{v\in C_{2}^{N}}|v+uH\rangle\right)\left(\sum_{v^{\prime}\in C_{2}^{N}}|v^{\prime}+xH\rangle\right)\left(\sum_{v^{\prime\prime}\in C_{2}^{N}}|v^{\prime\prime}+wH\rangle\right).$
This implies that $CCZ^{\otimes 2n}$ acts as identity on physical qubits. This completes the proof. ∎
The above theorem can be extended to the case when there are more than two controls. However, we skip that case here.
4 Conclusion
In this paper, we introduced a construction for generating binary CSS-T codes from any given binary CSS code. This construction enabled us to prove the existence of asymptotically good binary CSS-T codes, thereby resolving an open problem proposed in [32]. Moreover, we demonstrated that our construction preserves the LDPC property of the ingredient CSS code, leading to the existence of asymptotically good binary quantum LDPC codes that are also CSS-T. Finally, we extended our construction to support transversal finer $Z$ rotations and transversal $CCZ$, and analyzed their corresponding logical operators.
Our construction, as detailed in Section 2, relies on the use of a map $\phi$ satisfying the conditions of Proposition 1. In this paper, we chose $\phi$ to be the identity map as other permutation maps induce CSS-T codes with the same parameters. An interesting remaining question is to find other alternative maps $\phi$ satisfying the conditions of Proposition 1 and leading to CSS-T codes with better parameters or new applications (see also Remark 1).
Finally, another interesting follow-up would involve the possibility of using the proposed CSS-T construction in order to design magic state distillation protocols that may improve the overhead of existing protocols. Since most of the magic state distillation protocols are related to triorthogonal codes, it would be relevant to see how more general codes such as CSS-T codes may be used for such a task. While it is noteworthy that the transversal application of some of the non-Clifford gates considered in this paper results in the logical identity, such transformations have the potential to be used for magic state distillation [25, 26]. Furthermore, it would be relevant to study how the proposed codes can yield triorthogonal codes that have a more direct relationship with current magic state distillation protocols. Since our construction yields asymptotically good codes, successfully using those for magic state distillation tasks could result in making the magic state yield parameter arbitrarily small, which is still an open question for the qubit case.
Acknowledgment
This work was supported by the Spanish Ministry of Economy and Competitiveness through the MADDIE project (Grant No. PID2022-137099NBC44), by the Spanish Ministry of Science and Innovation through the project “Few-qubit quantum hardware, algorithms and codes, on photonic and solidstate systems” (PLEC2021-008251), by the Ministry of Economic Affairs and Digital Transformation of the Spanish Government through the QUANTUM ENIA project call - QUANTUM SPAIN project, and by the European Union through the Recovery, Transformation and Resilience Plan - NextGenerationEU within the framework of the Digital Spain 2026 Agenda, and by the grant ANR-21-CE39-0009-BARRACUDA from the French National Research Agency. J.E.M. is funded by the Spanish Ministry of Science, Innovation and Universities through a Jose Castillejo mobility grant for his stay at the Cavendish Laboratory of the University of Cambridge.